# Metric Oja depth, new statistical tool for estimating the most central objects

### v2

@@ 6 Conclusions
As shown in the simulations, MOD3 involves a significant time cost. By using partial U-statistics to randomly choose which elements to include in the triple sum in the sample version of (2), the speed of MOD3 computation could be enhanced. Another important point that could impact future results is the process of selecting and setting appropriate values for the hyper-parameters (take initial values as an example) used in the optimization algorithms.
Our timings in Section 4 did not take into account the computation of the interobject distances since this same step is required by all competing methods. In practice, the full runtime thus comprises of computing the distance matrix and the depths of the sample points. The specifics of the distance computation depend greatly on the metric space in question. For example, on the unit sphere in $\mathbb{R}^{p}$ with the arc length metric, forming the distance matrix is a $O(n^{2}p)$-operation, whereas on the manifold of $p\times p$ positive definite matrices with the Riemannian metric, this operation is of the order $O(n^{2}p^{3})$. The complexity of the depth computation depends only on $n$ and not on other aspects (such as dimensionality) of the sample, in a manner described in the beginning of Section 4.
The computation of MLD and MHD does not rely on the actual distances $d(x_{i},x_{j})$ but only on indicators of the form $\mathbb{I}\{d(x_{i},x_{j})\leq d(x_{i},x_{k})\}$, see Section 2. This means that MLD and MHD (a) are invariant to transformations that preserve these indicators, and (b) can be computed even if one knows only the binary values of these triple comparisons. The latter viewpoint for MLD was taken in [16] to find deepest objects based on binary dissimilarity data. Our MOD3, on the other hand, uses the actual values $d(x_{i},x_{j})$ in its computation, meaning that it cannot be computed based on indicators alone. This limits its usefulness in such dissimilarity scenarios but, at the same time, allows MOD3 to use additional information on the data, which shows as an improved performance in several of our experiments over the other methods. We also note that this behavior is perfectly in line with the original Oja depth, which also depends on the numerical values of the distances $\|x_{i}-x_{j}\|$.
Possible future research topics include: (a) Investigating why $D_{O2}$, despite not being a measure of depth, manages to emprically estimate central objects on the manifold of positive definite matrices and on the unit sphere. (b) Depth functions are typically expected to be maximized at a symmetric center of a distribution. However, unlike, e.g., the metric half-space depth, $D_{O3}$ does not have a natural companion concept of symmetry using which such a center could be defined. One possibility would be to use the quantities $A(x,X_{1},X_{2}):=\{d^{2}(x,X_{1})+d^{2}(x,X_{2})-d^{2}(X_{1},X_{2})\}/\{2d(x,X_{1})d(x,X_{2})\}$, which satisfy $A(x,X_{1},X_{2})\geq-1$ with equality if and only if $x$ lies in between of $X_{1}$ and $X_{2}$ [28, Theorem 1]. This suggests that if there exists an object $x_{0}$ such that
@@ Acknowledgments
The work of VZ was supported by the Finnish Doctoral Program Network in Artificial Intelligence, AI-DOC (decision number VN/3137/2024-OKM-6). The work of VZ and JV was supported by the Research Council of Finland (Grants 347501, 353769).

### v3

@@ 6 Conclusions
Possible future research topics include: (a) Selecting and setting appropriate values for the hyper-parameters (take initial values as an example) used in the optimization algorithms in Section 4.2. (b) Investigating why $D_{O2}$, despite not being a measure of depth, manages to emprically estimate central objects on the manifold of positive definite matrices and on the unit sphere. (c) Depth functions are typically expected to be maximized at a symmetric center of a distribution. However, unlike, e.g., the metric half-space depth, $D_{O3}$ does not have a natural companion concept of symmetry using which such a center could be defined. One possibility would be to use the quantities $A(x,X_{1},X_{2}):=\{d^{2}(x,X_{1})+d^{2}(x,X_{2})-d^{2}(X_{1},X_{2})\}/\{2d(x,X_{1})d(x,X_{2})\}$, which satisfy $A(x,X_{1},X_{2})\geq-1$ with equality if and only if $x$ lies in between of $X_{1}$ and $X_{2}$ [38, Theorem 1]. This suggests that if there exists an object $x_{0}$ such that
@@ Acknowledgments
The work of VZ was supported by the Finnish Doctoral Program Network in Artificial Intelligence, AI-DOC (decision number VN/3137/2024-OKM-6). The work of VZ and JV was supported by the Research Council of Finland (Grants 347501, 353769, 368494).